# Quantum Complexity in Constrained Many-Body Models: Scars, Fragmentation, and Chaos

### v1

@@ Abstract
Kinetic constraints in quantum many-body systems give rise to quantum states, whose behavior strongly depends on the choice of initial conditions. In recent years, these systems have drawn increasing interest because they provide insight into the mechanisms of thermalization and the situations where it can fail. In this work, we study a family of kinetically constrained models, including the celebrated Quantum Game of Life, from the perspective of quantum complexity, with a focus on entanglement, nonstabilizerness, and signatures of quantum chaos. By applying spectral diagnostics such as level statistics and spectral form factors, we demonstrate that these models—show robust chaotic behavior while also supporting Hilbert space fragmentation and quantum many-body scar states. Remarkably, we find that even certain symmetry-resolved fragmented sectors can themselves host scarred eigenstates, highlighting the unexpected coexistence of chaos, scars, and fragmentation within the same family of Hamiltonians. To better understand these fragmented subspaces, we further characterize them using their quantum resource generation ability. In particular, we demonstrate that characterization of entanglement and the ability to generate nonstabilizerness can be instrumental in distinguishing different dynamically disconnected sectors.
@@ I Introduction
Over the past decades, understanding thermalization in closed quantum systems has been a key question in quantum many-body physics. It is well known that integrable systems do not thermalize. This naturally leads to the question: do all non-integrable systems thermalize? One possible explanation for thermalization comes from the eigenstate thermalization hypothesis (ETH) [1, 2, 3, 4, 5, 6, 7], but there are notable exceptions [8, 9, 10]. Among the disorder-free spatially homogeneous quantum many-body models, kinetically constrained models (KCMs) often contain quantum states with unusual behavior that do not thermalize, commonly known as the quantum many-body scar (QMBS) states [11, 12, 13]. These scars offer a clear example of non-thermal behavior and show lower entanglement entropy. Recent experiments on a 51-atom quantum simulator provide direct evidence of quantum many-body scars [14]. See also Refs. [15, 16, 17] for other recent experimental implementations of QMBS. Moreover, recent studies have found another route to non-ergodicity through Hilbert space fragmentation (HSF) [18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. In these systems, the Hilbert space splits into dynamically disconnected sectors, which prevents full thermalization. Experimental examples of fragmentation have already been shown in platforms such as cold atoms and Rydberg arrays [21, 28, 29].
Physical systems that exhibit thermalization are of fundamental interest, yet they pose practical challenges for both analytical tractability and inherent computational complexity of simulating them. Non-interacting or integrable systems are relatively straightforward, as they can often be reduced to effective single-particle descriptions. Strongly interacting systems, however, resist such simplifications. This has motivated the search for quantifiers that capture how complex a quantum many-body system can be—both in preparing its states and in simulating its dynamics. In this context, entanglement [30] and non-stabilizerness [31] provide us with some insight regarding state-preparation complexity (see also [32, 33, 34]), while quantum chaos serves as an indicator of the difficulty of simulating the dynamics generated by a many-body Hamiltonian [35, 36]. In literature, there have been prior studies on such aspects. For example, a very recent work by Santra et al. [37] discusses quantum complexity in disordered models. Among the KCMs, Refs. [38, 39, 40] explore the quantum complexity properties of the PXP model and its variants [41], which serve as effective models of one-dimensional Rydberg atom quantum simulators. Nevertheless, there remain other classes of kinetically constrained models whose evolution rules can give rise to rich and interesting dynamical phenomena that are yet to be fully explored from this perspective.
@@
In our work, we investigate how different measures and markers of quantum complexity behave in one such family of kinetically constrained quantum many-body models in one-dimension (1D), where the dynamics of the quantum state at the ‘$i$’-th site depends on the total population of its neighbors, including both the nearest-neighbors (i1,i+1i-1,\,i+1) and the next-nearest-neighbors (i2,i+2i-2,\,i+2). A notable example is the Quantum Game of Life (QGL) [42, 43, 44], introduced as a quantum version of John Conway’s classical Game of Life [45]. Despite previous studies [46], it remains unclear whether this model exhibits genuine chaotic behavior. In our analysis, using the markers such as level statistics, and spectral form factor (SFF), we demonstrate that, along with the QGL, all other variants of KCMs considered in this work display clear signatures of quantum chaotic behavior when probed systematically. In particular, we explicitly show that symmetry plays a crucial role in revealing such features. However, for a subset of the models, mere symmetry resolution alone does not fully capture the spectral properties, as they exhibit Hilbert space fragmentation. In that case, to unveil the chaotic features, we need to study the dynamically disconnected sectors separately. Moreover, one of the fascinating results we uncover is the existence of KCMs for which the symmetry-resolved fragmented sectors further host quantum many-body scars. Therefore, by systematically varying the dynamical evolution constraints within these models, we can uncover the simultaneous emergence of chaotic behavior, quantum scars, and Hilbert space fragmentation within a single family of KCMs.
@@
Thereafter, we characterize these chaotic subspaces in terms of their potential for resource generation, focusing on entanglement and nonstabilizerness. Such a characterization is particularly crucial in the context of Hilbert space fragmentation. Unlike symmetry-resolved sectors, fragmented subspaces cannot be identified through conserved quantities. Instead, we employ complexity-based measures to probe them. In particular, we evaluate entanglement and nonstabilizerness as quantified by entanglement entropy [30] and stabilizer Rényi entropy [31], respectively, for the largest subspaces as well as several higher-dimensional fragmented ones, and compare their resource-generation capacity across dynamically disconnected sectors. Just as conserved quantities allow us to distinguish symmetry-resolved sectors, here we aim to differentiate fragmented subspaces through their complexity-driven resources.
Therefore, our findings highlight a wide range of complex physical phenomena that arise from a family of kinetically constrained models, even when they appear to be structurally similar at first glance. We hope that our findings will contribute to a deeper understanding of kinetically constrained models, with implications not only for fundamental physics but also for practical applications in quantum technologies such as some recent work that considers the application of these kinetically constrained models for quantum sensing [47, 48]. We stress here that in this work, we use the term quantum complexity specifically in the context of quantum state preparation and the simulation of its dynamics. In the broader literature, however, quantum complexity often refers to notions from quantum complexity theory and complexity classes [49], which lie beyond the scope of the present article.
@@
This work is organized as follows. In Section II, we introduce the family of kinetically constrained models that we examine in this study. Following this, in Section III, we discuss the effect of these kinetic constraints on the dimension of the active Hilbert space, including how they lead to Hilbert space fragmentation and their associated symmetries. In Section IV, we provide a detailed analysis of the quantum complexity aspects of these models, exploring entanglement, quantum chaos, and the behavior of stabilizer Rényi entropy. Sec. V further discusses the appearance of quantum many-body scar states in the dynamically disconnected fragmented sector and their robustness. The quantum complexity properties of the Hamiltonian $H_{\mathrm{Tot}}$ are discussed separately in Sec. VI, highlighting key aspects that differ from those observed in other models. Finally, in Section VII, we present a discussion on the future outlook of our work.
@@ II Model Hamiltonians
Before we present the details of our analysis, in this section, we define the set of quantum many-body Hamiltonians that we have considered in our work. One such primary candidate is the Hamiltonian $H_{\mathrm{Tot}}$, acting on a 1D spin-1/2 chain, which we define below.
$H_{\mathrm{Tot}}$ $=\sum_{i=1}^{L}\sigma^{x}_{i}(\mathcal{N}_{i}^{(0)}+\mathcal{N}_{i}^{(1)}+\mathcal{N}_{i}^{(2)}+\mathcal{N}_{i}^{(3)}),$
$=H^{\mathcal{N}(0)}+H^{\mathcal{N}(1)}+H^{\mathcal{N}(2)}+H^{\mathcal{N}(3)},$ (1)
where $\sigma^{x}_{i}$ is the Pauli $x$ operator at site $i$, and the Hamiltonians $H^{\mathcal{N}(k)}$ described above each represent a hierarchical level of kinetic constraints that govern the dynamics at site $i$, as determined by the operator $\mathcal{N}_{i}^{(k)}$. Specifically, the $\mathcal{N}_{i}^{(k)}$’s are operators defined on four adjacent sites of $i$ (left and right nearest neighbors i1,i+1i-1,i+1 and left and right next-nearest neighbors i2,i+2i-2,i+2) that ensure the evolution criterion is met: the quantum state at site $i$ will flip provided the sum of the population of its four adjacent sites equals $k$. Mathematically, we can define them as follows:
@@
with $n_{k}=|0\rangle\langle 0|_{k}$ and $\bar{n}_{k}=|1\rangle\langle 1|_{k}$.
In our work, along with the quantum properties of the Hamiltonian defined in Eq. (1) and individual $H^{\mathcal{N}(k)}$’s, we also consider quantum Hamiltonians that arise from various combinations of $H^{\mathcal{N}(k)}$. In what follows, we briefly introduce them for future reference.
(a) Perturbed PPXPP model: The first case we consider in our work is a combination of $H^{\mathcal{N}(0)}$ (unconstrained PPXPP) and $H^{\mathcal{N}(1)}$. The resulting Hamiltonian is denoted by
$H^{\text{Pert}}_{\text{PPXPP}}=H^{\mathcal{N}(0)}+\delta\cdot H^{\mathcal{N}(1)}=\sum_{i=1}^{L}\sigma^{x}_{i}(\mathcal{N}_{i}^{(0)}+\delta\cdot\mathcal{N}_{i}^{(1)}),$ (3)
where $\delta$ is a constant. Note that, unlike the conventional PPXPP model, here, $H^{\mathcal{N}(0)}$ does not include the Rydberg constraint, which restricts the basis to have configurations where any site in state $|1\rangle$ is always surrounded by at least two zeros on both its left and right, i.e., of the form $|\dots 00100\dots\rangle$. In Sec. V, we discuss how this model lead to the remarkable coexistence of quantum many-body scars and Hilbert space fragmentation within it.
@@
In the literature, there have been previous proposals for implementation of quantum kinetically constrained models that belong to the paradigm of quantum cellular automata [52] with Rydberg atoms [53]. The main idea is to use programmable multifrequency excitation and depumping of Rydberg states to create conditional interactions that are similar to classical cellular automata. In our case, we conjecture that the Quantum Game of Life, a specific form of cellular automaton and other related models could be realized using similar experimental approaches.
@@ III Hamiltonian-Induced Hilbert Space Structure
III.1 Active Hilbert space dimension
One such direct consequence of the kinetic constraints is the annihilation of certain basis states. As a result, the system is unable to fully explore the entire Hilbert space, since some regions become dynamically inaccessible. The effective Hilbert space dimension in this case depends on the type of constraints and the combination of the Hamiltonians considered. In Table 1, we list them down for different $H^{\mathcal{N}(k)}$’s as well as $H^{\text{QGL}}$ and $H^{\text{Pert}}_{\text{PPXPP}}$.
@@ III.1 Active Hilbert space dimension
Table 1: Scaling of the effective Hilbert space dimensions ($d$) with system size $L$ of kinetically constrained Hamiltonians $H^{\mathcal{N}(k)}$, for $k=0,1,2,3$ and the Hamiltonians constructed as their combinations, $H^{\text{Pert}}_{\text{PPXPP}}$ and $H^{\text{QGL}}$.
@@ III.2 Symmetry resolutions
(a) Translational symmetry: With periodic boundary conditions, all the Hamiltonians $\{H^{\mathcal{N}(k)},H^{\text{QGL}},H^{\text{Pert}}_{\text{PPXPP}}\}$ exhibit translation symmetry. Let $\hat{\mathcal{T}}$ be the translation operator which shifts the $i$’th site to the $(i+1)$’th site. This operator commutes with all the Hamiltonians. The eigenvalue of this operator is $e^{ik}$, where $k$ is the momentum of that eigenstate. In our case, we primarily focus on the translationally invariant zero-momentum ($k=0$) subspace.
@@
(d) Spin-flip or global $\mathbb{Z}_{2}$ symmetry: The final symmetry we will discuss here is the spin-flip or also known as global $\mathbb{Z}_{2}$ symmetry exhibited by $H^{\mathcal{N}^{(2)}}$. This symmetry commutes with the spin-flip operator $\hat{\kappa}=\prod_{i}\sigma^{x}_{i}$. The operator has eigenvalues of $\kappa=\pm 1$, leading to symmetric and antisymmetric spin-flip sectors. For the rest of the Hamiltonians, the effect of this local spin-flip operation results in

### v2

# Quantum Complexity in Rule-Based Constrained Many-Body Models: Scars, Fragmentation, and Chaos

@@ Abstract
Kinetic constraints in quantum many-body systems strongly restrict the accessible Hilbert space, giving rise to highly nontrivial dynamical behavior. In recent years, such systems have attracted growing interest as they provide insight into mechanisms of thermalization and into regimes where thermalization fails. In this work, we study a family of rule-based kinetically constrained models, including the celebrated Quantum Game of Life, from the perspective of quantum complexity, with a focus on entanglement, nonstabilizerness, and quantum chaos. Using spectral diagnostics such as level statistics and spectral form factors, we show that these models exhibit robust chaotic behavior while simultaneously supporting both strong and weak Hilbert-space fragmentation and quantum many-body scar states. To further elucidate the structure of these fragmented subspaces, we characterize them through their ability to generate quantum resources. In particular, we show that resource-generation capacity does not necessarily correlate with the dimensionality of a fragmented sector, and that entanglement structure and the ability to generate nonstabilizerness provide effective diagnostics for distinguishing dynamically disconnected sectors. Our work therefore explores kinetically constrained models in a general framework that is not restricted to Rydberg blockade–based constraints alone.
@@ I Introduction
Over the past decades, understanding thermalization in closed quantum systems has been a key question in quantum many-body physics. It is well known that integrable systems do not thermalize in the conventional sense, and instead relax to generalized Gibbs ensembles. This naturally leads to the question: do all non-integrable systems thermalize? One possible explanation for thermalization comes from the eigenstate thermalization hypothesis (ETH) [1, 2, 3, 4, 5, 6, 7], but there are notable exceptions [8, 9, 10, 11]. Among the disorder-free spatially homogeneous quantum many-body models, kinetically constrained models (KCMs) often contain quantum states with unusual behavior that do not thermalize, commonly known as the quantum many-body scar (QMBS) states [12, 13, 14]. These scars offer a clear example of non-thermal behavior and show lower entanglement entropy. Recent experiments on a 51-atom quantum simulator provide direct evidence of quantum many-body scars [15]. See also Refs. [16, 17, 18] for other recent experimental implementations of QMBS. Moreover, recent studies have found another route to non-ergodicity through Hilbert space fragmentation (HSF) [19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. In these systems, the Hilbert space splits into dynamically disconnected sectors, which prevents full thermalization. Experimental examples of fragmentation have already been shown in platforms such as cold atoms and Rydberg arrays [22, 29, 30].
In recent years, much of the interest in studying kinetically constrained models has been driven by their experimental realization in Rydberg atom arrays operating in the blockade regime. This experimental accessibility has opened avenues to explore a broader range of physical phenomena. For instance, while the PXP model is known to host quantum many-body scar states, it does not exhibit Hilbert-space fragmentation; such fragmentation has been shown to emerge in Floquet generalizations of the PXP model [20], exhibiting coexistence of both scar and fragmentation. More recently, related forms of constrained dynamics have also been explored in the antiblockade regime of Rydberg systems, where facilitation mechanisms lead to distinct dynamical rules and constraint structures [31].
In parallel, recent years have also witnessed growing interest in kinetically constrained quantum dynamics formulated in terms of rule-based evolution, where local updates depend on the population of neighboring sites rather than on a fixed symmetric projector structure. A prominent example is the Quantum Game of Life (QGL) [32, 33, 34], introduced as a quantum analogue of John Conway’s classical Game of Life [35], in which the dynamics at a given site is governed by the total occupation of its surrounding sites.
Motivated by the rule-based perspective exemplified by the Quantum Game of Life, and using symmetric projector-based PXP-type models as reference points, we focus in this work on a family of one-dimensional kinetically constrained quantum many-body models in which the dynamics at site $i$ depends on the total population of its neighboring sites, including both nearest neighbors (i1,i+1)(i-1,i+1) and next-nearest neighbors (i2,i+2)(i-2,i+2). When expressed in operator form, such population-dependent dynamical rules naturally give rise to sum of asymmetric projectors, in contrast to the symmetric constraints appearing in standard PXP- or PPXPP-type models, where identical projectors act on both sides of a site. In other words, this provides a route to explore kinetic constraints beyond strict blockade conditions.
@@
Despite earlier investigations [36], it remains unclear whether the Quantum Game of Life and related kinetically constrained models exhibit genuine quantum chaotic behavior. Using standard chaos diagnostics such as level-spacing statistics and the spectral form factor (SFF), we demonstrate that, beyond the QGL, all variants of kinetically constrained models considered in this work display clear signatures of quantum chaos when analyzed systematically. A key observation is that symmetry resolution plays a crucial role in revealing these features. However, for a subset of the models, resolving symmetries alone is insufficient: the spectrum further exhibits Hilbert space fragmentation, requiring an analysis at the level of dynamically disconnected sectors. Only upon separately probing these sectors do chaotic signatures become apparent. Remarkably, we find that in certain models the symmetry-resolved fragmented sectors themselves host quantum many-body scars [37]. Moreover, the HSF patterns observed among different members of the KCM family are markedly distinct, exhibiting strong, weak, and moderate fragmentation. Thus, by systematically tuning the dynamical constraints, we uncover the simultaneous emergence of quantum chaos, Hilbert space fragmentation, and many-body scarring within a single family of kinetically constrained models.
@@
We then characterize these chaotic subspaces from the perspective of quantum resource generation, focusing on entanglement and non-stabilizerness. In this context, entanglement and non-stabilizerness provide us with some insight regarding state-preparation complexity (see also [38, 39, 40, 41]). In literature, there have been prior studies on such aspects. For example, a very recent work by Santra et al. [42] discusses quantum complexity in disordered models. Among the KCMs, Refs. [43, 44, 45] explore the quantum complexity properties of the PXP model and its variants [46, 47]. In our case, this characterization is particularly important in the presence of Hilbert space fragmentation: unlike symmetry-resolved sectors, fragmented subspaces cannot be identified through conventional conserved quantities. Instead, we probe them using complexity-based measures. Specifically, we evaluate the entanglement entropy [48] and the stabilizer Rényi entropy [49] for the largest connected component as well as several higher-dimensional fragmented subspaces, and compare their capacity for resource generation across dynamically disconnected sectors. In this way, just as conserved quantities distinguish symmetry sectors, complexity measures allow us to differentiate fragmented subspaces through their static and dynamical resources.
Therefore, our findings reveal a broad range of complex physical phenomena that emerge across a family of kinetically constrained quantum many-body models, even when these models appear structurally similar at first glance. By showing that quantum many-body scars, Hilbert-space fragmentation, and slow dynamics persist in systems governed by population-dependent rules rather than symmetric projector-based constraints, we demonstrate that these phenomena are not fine-tuned to a specific Hamiltonian form (comprising of Rydberg blockade). Instead, they arise more generally from the underlying constraint logic, thereby broadening the theoretical framework of kinetically constrained quantum dynamics. We reiterate that, although several works have addressed quantum cellular automata and other rule-based models, a systematic and unified characterization of different notions of quantum complexity in such systems has so far been lacking that we addressed in our work. Beyond their fundamental significance, our results may also be relevant for emerging quantum technologies, particularly in light of recent proposals that exploit kinetically constrained models for applications such as quantum sensing [50, 51]. We stress here that in this work, we use the term quantum complexity specifically in the context of quantum state preparation and the simulation of its dynamics. In the broader literature, however, quantum complexity often refers to notions from quantum complexity theory and complexity classes [52], which lie beyond the scope of the present article.
@@
This work is organized as follows. In Section II, we introduce the family of kinetically constrained models that we examine in this study. Following this, in Section III, we discuss the effect of these kinetic constraints on the dimension of the effective Hilbert space, including how they lead to Hilbert space fragmentation and their associated symmetries. In Section IV, we provide a detailed analysis of the quantum complexity aspects of these models, exploring entanglement, stabilizer Rényi entropy and quantum chaos. Sec. V further discusses the appearance of quantum many-body scar states in the dynamically disconnected fragmented sector and their robustness. The quantum complexity properties of the Hamiltonian $H^{\mathrm{Tot}}$ are discussed separately in Sec. VI, highlighting key aspects that differ from those observed in other models. Finally, in Section VII, we present a discussion on the future outlook of our work.
@@ II Model Hamiltonians
Before we present the details of our analysis, in this section, we define the set of rule-based quantum many-body Hamiltonians that we have considered in our work. One such primary candidate is the Hamiltonian $H^{\mathrm{Tot}}$, acting on a 1D spin-1/2 chain, which we define below.
$H^{\mathrm{Tot}}$ $=\sum_{i=1}^{L}\sigma^{x}_{i}(\mathcal{N}_{i}^{(0)}+\mathcal{N}_{i}^{(1)}+\mathcal{N}_{i}^{(2)}+\mathcal{N}_{i}^{(3)}),$
$=H^{\mathcal{N}(0)}+H^{\mathcal{N}(1)}+H^{\mathcal{N}(2)}+H^{\mathcal{N}(3)}.$ (1)
We use periodic boundary condition (PBC) throughout the paper. Here $\sigma^{x}_{i}$ is the Pauli $x$ operator at site $i$, and the Hamiltonians $H^{\mathcal{N}(k)}$ described above each represents a hierarchical level of kinetic constraints that govern the dynamics at site $i$, as determined by the operator $\mathcal{N}_{i}^{(k)}$. Specifically, the $\mathcal{N}_{i}^{(k)}$’s are operators defined on four adjacent sites of $i$ (left and right nearest neighbors i1,i+1i-1,i+1 and left and right next-nearest neighbors i2,i+2i-2,i+2) that ensure the evolution criterion is met: the quantum state at site $i$ will flip provided the sum of the population of its four adjacent sites equals $k$. Mathematically, we can define them as follows:
@@
with $n_{k}=|0\rangle\langle 0|_{k}$ and $\bar{n}_{k}=|1\rangle\langle 1|_{k}$. From the above expression we can see $H^{\mathcal{N}(0)}$ is unconstrained PPXPP model that unlike the conventional PPXPP model, does not include the Rydberg constraint, which restricts the basis to have configurations where any site in state $|1\rangle$ is always surrounded by at least two zeros on both its left and right, i.e., of the form $|\dots 00100\dots\rangle$. Moreover, Eq. (2) explicitly shows that each $H^{\mathcal{N}(k)}$ consists of terms involving asymmetric projectors.
In our work, along with the quantum properties of the Hamiltonian defined in Eq. (1) and individual $H^{\mathcal{N}(k)}$’s, we also consider quantum Hamiltonians that arise from various combinations of different $H^{\mathcal{N}(k)}$. In what follows, we briefly introduce them for future reference.
(a) Perturbed PPXPP model: The first case we consider in our work is a combination of $H^{\mathcal{N}(0)}$ (unconstrained PPXPP) and $H^{\mathcal{N}(k)}$, where $k=1$ and 2. The resulting Hamiltonian is denoted by
$H^{\text{Pert}}_{\text{PPXPP}}(k)=H^{\mathcal{N}(0)}+\delta\cdot H^{\mathcal{N}(k)}=\sum_{i=1}^{L}\sigma^{x}_{i}(\mathcal{N}_{i}^{(0)}+\delta\cdot\mathcal{N}_{i}^{(k)}),$ (3)
where $\delta$ is a constant. In Sec. V, we discuss how this model lead to the remarkable coexistence of quantum many-body scars and Hilbert space fragmentation within it.
@@
In the literature, there have been previous proposals for implementation of quantum kinetically constrained models that belong to the paradigm of quantum cellular automata [55] with Rydberg atoms [56]. The main idea is to use programmable multifrequency excitation and depumping of Rydberg states to create conditional interactions that are similar to classical cellular automata. Importantly, the essential ingredient in such realizations is not a specific blockade-induced Hamiltonian, but the ability to implement local, rule-dependent evolution conditioned on the surrounding population. In our case, we conjecture that the rule-based KCMs including the Quantum Game of Life, could also be realized using similar experimental approaches.
@@ III Hamiltonian-Induced Hilbert Space Structure
III.1 Effective Hilbert space dimension
One such direct consequence of the kinetic constraints is the annihilation of certain basis states. As a result, the system is unable to fully explore the entire Hilbert space, since some regions become dynamically inaccessible. The effective Hilbert space dimension (excluding the trivially annihilated states) in this case depends on the type of constraints and the combination of the Hamiltonians considered. In Table 1, we list them down for different $H^{\mathcal{N}(k)}$’s as well as $H^{\text{QGL}}$.
@@ III.1 Effective Hilbert space dimension
Table 1: Scaling of the effective Hilbert space dimensions ($d$) with system size $L$ of kinetically constrained Hamiltonians $H^{\mathcal{N}(k)}$, for $k=0,1,2,3$ and $H^{\text{QGL}}$
@@ III.2 Symmetry resolutions
(a) Translational symmetry: With periodic boundary conditions, all the Hamiltonians $\{H^{\mathcal{N}(k)},H^{\text{QGL}},H^{\text{Pert}}_{\text{PPXPP}}(k)\}$ exhibit translation symmetry. Let $\hat{\mathcal{T}}$ be the translation operator which shifts the $i$’th site to the $(i+1)$’th site. This operator commutes with all the Hamiltonians. The eigenvalue of this operator is $e^{ik}$, where $k$ is the momentum of that eigenstate. In our case, we primarily focus on the translationally invariant zero-momentum ($k=0$) subspace.
@@
(d) Spin-flip or global $\mathbb{Z}_{2}$ symmetry: The final symmetry we will discuss here is the spin-flip or also known as global $\mathbb{Z}_{2}$ symmetry exhibited by $H^{\mathcal{N}(2)}$. This Hamiltonian commutes with the spin-flip operator $\hat{\kappa}=\prod_{i}\sigma^{x}_{i}$. The operator has eigenvalues of $\kappa=\pm 1$, leading to symmetric and antisymmetric spin-flip sectors. For the rest of the Hamiltonians, the effect of this local spin-flip operation results in
@@
Figure 1: Krylov subspaces and their scaling in fragmented model $H^{\mathcal{N}(1)}$. In (a) and (b), we illustrate how the number of Krylov subspaces, $\mathcal{N}_{L}$, and the ratio $d_{L}/\mathcal{D}_{L}$ behaves with system size $L$, respectively. The maximum system size considered is $L=22$. Both figures clearly indicate the presence of strong Hilbert space fragmentation (HSF). In contrast, $H^{\mathcal{N}(2)}$ (shown in Appendix B) displays a slower decay of $d_{L}/\mathcal{D}_{L}$ with increasing $L$.